# Towards Embodied Agentic AI: Review and Classification of LLM- and VLM-Driven Robot Autonomy and Interaction

### v2

@@ III Approaches to LLMs/VLMs/VLAs and AI Agents in Robotics
To that end, we propose the existence of, currently, four main directions in which foundation models are integrated into robotic systems: (i) protocol integration, where the focus is on using the model as a translator between protocols (e.g., user input to predefined toolset); (ii) interface integration, where the focus is to provide an interactive approach connecting user, robotic system and the environment; (iii) orchestration-oriented integration, where the model is in charge of managing resources, tools or subsystems; and (iv) direct or embedded integration, where the model acts as a perception or control policy, either end-to-end, or as a specific subsystem. These approaches are illustrated in Fig. 1. We also classify a selection of representative works based on their integration approach in Table II.
@@ III-A Protocol Integration
For example, ros2ai [22] does this by translating a generic user prompt into one specific ROS 2 command line interface tool call. The interaction is limited, although possible, and the process is mostly unidirectional.
More recently, with the advent of the MCP protocol as the de-facto standard for tool use in LLMs/VLMs, multiple authors have proposed technical integrations [28, 25]. The available tooling can span multiple domains: interaction with graphical interfaces (e.g., launching a simulator), the previously mentioned execution of ROS 2 command line interface commands, or the connectivity with ROS communication paradigms (publishing / subscribing to specific topics, calling services, or executing actions. Indeed, the MCP protocol maps well with, e.g., the service and action interfaces in ROS.
A notable related approach is code-as-policies [6], where the LLM outputs Python code to perform the task (using a predefined API for sensorimotor actions). The code is then executed in a loop. Essentially the LLM “programs” the robot on the fly.
@@ III-B Interface Integration
The most recent and arguably complete open-source agentic system to date is OpenMind [16], a commercial effort that has been demonstrated in humanoid and quadruped robots, among others, and is defining a new AI-centric ”operating system” for intelligent robots.
@@ III-D Embedded or Direct Integration
This category encompasses approaches where the LLM (or a multimodal variant) directly produces either (i) actions for the robot in an end-to-end manner, or (ii) a specific output (e.g., to be used as a perception module). These are often referred to as robotic foundation models, in their end-to-end variant. These approaches are also well covered in existing review papers [1, 2, 3, 4]. Therefore, we only include in this survey those that are most relevant from a historical perspective, or have potential to advance towards more agentic systems. Those works are included in Table II.
@@ III-E Agentic Frameworks
Figure 3: Taxonomy of Robotic Agents grouped by role and functionality.
@@ IV Roles and Architectures of Robotic Agents with LLMs
As we surveyed recent LLM-integrated robotic systems, it became clear that many differ not in the tasks they perform, but in how they structure decision-making, control, and modularity. To capture these architectural distinctions, we group the agents according to their functional design rather than their application domain. This perspective enables a clearer comparison of how different systems reason, sequence actions, invoke tools or skills, and interface with low-level controllers. An overview of these agent types is provided in Figure 3.
@@ IV-1 Planner Agents
A foundational example is Google’s SayCan [34], where the LLM (PaLM) generates possible next actions (e.g., “pick up can” or “move to kitchen”), and a learned value function (affordance model) evaluates each option based on the robot’s current state. The most feasible action is then executed by the appropriate skill module. Similarly, SELP [35] uses an LLM to generate symbolic task plans, followed by structured safety and efficiency filtering before execution. ConceptAgent [36] extends this idea further by incorporating a symbolic planner and a precondition-grounding module, enabling robust task decomposition and dynamic re-planning based on environment feedback.
@@ IV-2 Orchestration Agents
For example, AutoRT [27] deploys an LLM to orchestrate a fleet of real-world mobile manipulators. The system uses prompt-based reasoning to map natural language instructions to specific skill invocations (e.g., grasping, opening, navigating), while integrating filtering layers to ensure safety and feasibility. Similarly, LABOR Agent [38] enables bimanual robotic manipulation by selecting among hundreds of pretrained skills. The LLM determines which skill or skill chain should be used given a complex manipulation instruction, using grounding modules to connect symbolic task descriptions to real-world motor behaviors.
In our categorization, orchestration agents primarily refer to systems that handle high-level task planning and assignment, either by coordinating multiple internal agents within a single robot [37], or by distributing tasks across multiple embodied agents (i.e., multiple robots) [39]. While they rely on predefined capabilities, the orchestration logic remains flexible, enabling adaptive responses to novel combinations of goals, environments, or agent configurations.
@@ IV-3 Task-Specific Agents
Task-specific agents are designed to solve narrowly defined problems, which used to rely on specialized training data or domain-specific policies. Recent work, however, explores how large language and vision-language models can enhance task-specific performance by enabling zero-shot reasoning or dynamic planning for particular objectives. For instance, LGX enables an embodied agent to find novel, uniquely described objects (e.g., ”cat-shaped mug”) using language-guided exploration and vision-language grounding without task-specific training [42]. Similarly, NavGPT uses explicit reasoning over visual and linguistic cues to follow navigation instructions, demonstrating that language models can guide task-specific behavior without prior tuning on navigation data [41]. ARNA leverages an LVLM to dynamically orchestrate perception and control modules for general-purpose navigation, but it can also be framed as a task-specific agent when executing goal-directed exploration in unseen environments [30]. Similarly, BUMBLE demonstrates a highly capable agent for manipulation and navigation by integrating an LLM with a world model and structured memory, enabling it to follow user instructions across tasks.
@@ IV-4 Model-Centric Agents
Unlike traditional VLA models that often require separate components for perception, reasoning, and action, each trained or tuned independently, model-centric agents adopt a unified architecture where a single model is responsible for grounding multimodal inputs (e.g., images, language, proprioception) and directly producing action outputs. Examples include LEO [43], which leverages a decoder-only large language model to integrate 2D egocentric vision, 3D point clouds, and text for both instruction following and physical interaction in 3D environments; RoboCat [44], which uses a goal-conditioned decision transformer to generalize across robotic embodiments and tasks through large-scale training and self-improvement; and RoboAgent [45], which achieves high data efficiency and broad task generalization through semantic augmentation and action chunking within a unified policy model.
@@ IV-5 Generalist Agents
For example, Voyager [32] is an open-ended generalist agent that autonomously explores new tasks in Minecraft by generating its own tools (as Python functions), evaluating them, and storing them for future use—effectively building a lifelong, self-curated skill library. Similarly, Code as Policies uses an LLM to generate executable Python policies that map observations directly to robot actions. ODYSSEY [33] likewise operates in an open-world setting, using an LLM to reason about tasks and select from a rich library of skills grounded in the environment. In the real world, RoboGPT [46] demonstrates generalist behavior by interpreting language instructions and invoking appropriate pretrained skills for manipulation and navigation, without needing task-specific retraining.
@@ IV-6 Generalist Systemic Agents
V Conclusion

### v3

@@ III Approaches to LLMs/VLMs/VLAs and AI Agents in Robotics
To that end, we propose the existence of, currently, four main directions in which foundation models are integrated into robotic systems: (i) protocol integration, where the focus is on using the model as a translator between protocols (e.g., user input to predefined toolset); (ii) interface integration, where the focus is to provide an interactive approach connecting user, robotic system and the environment; (iii) orchestration-oriented integration, where the model is in charge of managing resources, tools or subsystems; and (iv) direct or embedded integration, where the model acts as a perception or control policy, either end-to-end, or as a specific subsystem. These approaches are illustrated in Fig. LABEL:fig:concept. Table I also provides a summary of the principal aspects and differences within this categorization. In addition, we classify a selection of representative works based on their integration approach in Table II.
@@ III-A Protocol Integration
For example, ros2ai [ros2ai] does this by translating a generic user prompt into one specific ROS 2 command line interface tool call. The interaction is limited, although possible, and the process is mostly unidirectional (e.g., the user is presented with the response of the ROS command).
More recently, with the advent of the MCP protocol as the de-facto standard for tool use in LLMs/VLMs, multiple authors have proposed technical integrations [rosmcp, rosmcpserver]. The available tooling can span multiple domains: interaction with graphical interfaces (e.g., launching a simulator), the previously mentioned execution of ROS 2 command line interface commands, or the connectivity with ROS communication paradigms (publishing/subscribing) to specific topics, calling services, or executing actions. Indeed, the MCP protocol maps well with, e.g., the service and action interfaces in ROS.
A notable related approach is code-as-policies [liang2022code], where the LLM outputs Python code to perform the task (using a predefined API for sensorimotor actions). The code is then executed in a loop. Essentially, the LLM “programs” the robot on the fly.
@@ III-B Interface Integration
The most recent and arguably complete open-source agentic system to date is OpenMind [openmind], a commercial effort that has been demonstrated in humanoid and quadruped robots, among others, and is defining a new AI-centric “operating system” for intelligent robots.
@@ III-D Embedded or Direct Integration
This category encompasses approaches where the LLM (or a multimodal variant) directly produces either (i) actions for the robot in an end-to-end manner, or (ii) a specific output (e.g., to be used as a perception module). These are often referred to as robotic foundation models, in their end-to-end variant. These approaches are also well covered in existing review papers [firoozi2025foundation, urain2024deep, zeng2023large, xiao2025robot]. Therefore, we only include in this survey those that are most relevant from a historical perspective, or have potential to advance towards more agentic systems. Those works are included in Table II.
@@ III-E Agentic Frameworks
Robotic Agents
[…]
Figure 3: Classification of existing works by primary role and functionality.
@@ IV Roles and Architectures of Robotic Agents with LLMs
As we surveyed recent LLM-integrated robotic systems, it became clear that many differ not in the tasks they perform, but in how they structure decision-making, control, and modularity. To capture these architectural distinctions, we group the agents according to their functional design rather than their application domain. This perspective enables a clearer comparison of how different systems reason, sequence actions, invoke tools or skills, and interface with low-level controllers. While the boundaries between categories are often blurred, since many systems integrate multiple capabilities, the classification in Figure 3 highlights each work’s primary role and functionality.
@@ IV-1 Planner Agents
A foundational example is Google’s SayCan [ahn2022can], where the LLM (PaLM) generates possible next actions (e.g., “pick up can” or “move to kitchen”), and a learned value function (affordance model) evaluates each option based on the robot’s current state. The most feasible action is then executed by the appropriate skill module. Similarly, SELP [wu2024selp] uses an LLM to generate symbolic task plans, followed by structured safety and efficiency filtering before execution. ConceptAgent [rivera2024conceptagent] extends this idea further by incorporating a symbolic planner and a precondition-grounding module, enabling robust task decomposition and dynamic re-planning based on environment feedback.
@@ IV-2 Orchestration Agents
For example, AutoRT [ahn2024autort] deploys an LLM to orchestrate a fleet of real-world mobile manipulators. The system uses prompt-based reasoning to map natural language instructions to specific skill invocations (e.g., grasping, opening, navigating), while integrating filtering layers to ensure safety and feasibility. Similarly, LABOR Agent [chu2024large] enables bimanual robotic manipulation by selecting among hundreds of pretrained skills. The LLM determines which skill or skill chain should be used given a complex manipulation instruction, using grounding modules to connect symbolic task descriptions to real-world motor behaviors.
In our categorization, orchestration agents primarily refer to systems that handle high-level task planning and assignment, either by coordinating multiple internal agents within a single robot [glocker2025llm], or by distributing tasks across multiple embodied agents (i.e., multiple robots) [kannan2024smart]. While they rely on predefined capabilities, the orchestration logic remains flexible, enabling adaptive responses to novel combinations of goals, environments, or agent configurations.
@@ IV-3 Task-Specific Agents
Task-specific agents are designed to solve narrowly defined problems, which used to rely on specialized training data or domain-specific policies. Recent work, however, explores how large language and vision-language models can enhance task-specific performance by enabling zero-shot reasoning or dynamic planning for particular objectives. For instance, LGX enables an embodied agent to find novel, uniquely described objects (e.g., ”cat-shaped mug”) using language-guided exploration and vision-language grounding without task-specific training [dorbala2023can]. Similarly, NavGPT uses explicit reasoning over visual and linguistic cues to follow navigation instructions, demonstrating that language models can guide task-specific behavior without prior tuning on navigation data [zhou2024navgpt]. ARNA leverages an LVLM to dynamically orchestrate perception and control modules for general-purpose navigation, but it can also be framed as a task-specific agent when executing goal-directed exploration in unseen environments [lange2025general]. Similarly, BUMBLE demonstrates a highly capable agent for manipulation and navigation by integrating an LLM with a world model and structured memory, enabling it to follow user instructions across tasks.
@@ IV-4 Model-Centric Agents
Unlike traditional VLA models that often require separate components for perception, reasoning, and action, each trained or tuned independently, model-centric agents adopt a unified architecture where a single model is responsible for grounding multimodal inputs (e.g., images, language, proprioception) and directly producing action outputs. Examples include LEO [huang2023embodied], which leverages a decoder-only large language model to integrate 2D egocentric vision, 3D point clouds, and text for both instruction following and physical interaction in 3D environments; RoboCat [bousmalis2023robocat], which uses a goal-conditioned decision transformer to generalize across robotic embodiments and tasks through large-scale training and self-improvement; and RoboAgent [bharadhwaj2024roboagent], which achieves high data efficiency and broad task generalization through semantic augmentation and action chunking within a unified policy model.
@@ IV-5 Generalist Agents
For example, Voyager [wang2023voyager] is an open-ended generalist agent that autonomously explores new tasks in Minecraft by generating its own tools (as Python functions), evaluating them, and storing them for future use—effectively building a lifelong, self-curated skill library. Similarly, Code as Policies uses an LLM to generate executable Python policies that map observations directly to robot actions. ODYSSEY [liu2024odyssey] likewise operates in an open-world setting, using an LLM to reason about tasks and select from a rich library of skills grounded in the environment. In the real world, RoboGPT [chen2025robogpt] demonstrates generalist behavior by interpreting language instructions and invoking appropriate pretrained skills for manipulation and navigation, without needing task-specific retraining.
@@ IV-6 Generalist Systemic Agents
V Toolkits for Agentic Physical AI
Early agent systems in artificial intelligence, such as symbolic planners, belief–desire–intention (BDI) architectures, and the SOAR cognitive framework, laid the groundwork for connecting reasoning with action. The advent of deep learning shifted the focus to data-driven methods, and more recently, LLMs have enabled agents to integrate reasoning, perception, and action in naturalistic ways. Toolkits for agentic physical AI now aim to close the loop between high-level reasoning and embodied control, making it possible to deploy agents in robotics and other physical domains. This section surveys the main components of this emerging ecosystem.
V-A Large Language Models and Providers
State-of-the-art LLMs form the backbone of agentic systems. Closed-source providers such as OpenAI (GPT-4o), Anthropic (Claude 3.5), and Google DeepMind (Gemini 1.5) offer advanced reasoning and multimodal capabilities, along with integrated tool use. Open-source efforts include Meta (Llama 3.1), Mistral AI (Mixtral MoE), and xAI (Grok). These models differ in openness, context length, and function-calling capabilities, but all increasingly support structured integration with external systems.
V-B Tool Calling
While impressive, LLMs originally operated only through isolated textual interfaces limited to their training data, without the ability to directly access real-time information, external tools, or APIs. To bridge this gap, one of the first major steps was function calling, introduced by OpenAI, where models were fine-tuned to recognize when a function should be invoked and respond with structured JSON that matched the function’s signature [buadicua2025contemporary]. Over time, this approach evolved into what is now broadly known as tool calling (renamed from function calling), which is supported by most modern LLMs. However, implementations differ across vendors in how JSON structures are defined and handled directly by LLMs. The Model Context Protocol (MCP) builds on this idea by introducing an open, interoperable standard that enables consistent and reliable communication between clients (LLM-powered applications) and servers (e.g., tools). Unlike vendor-specific tool calling, MCP ensures that hosts can access tools, resources, and real-time context updates in a standardized way without constant custom integration work.
In the MCP, there are three key components: the MCP Server, MCP Client, and MCP Host. The MCP Server is a program that provides context in the form of tools (callable functions or actions), resources (structured data or knowledge), and prompts (predefined templates). A server can define multiple of these and communicate them using the JSON-RPC 2.0 protocol, either locally through stdio transport or remotely through streaming HTTP. The MCP Client is a protocol-level component that connects to exactly one server and is responsible for handling this communication, passing the context back to the host. Finally, the MCP Host is any AI application that integrates one or more MCP Clients to extend its capabilities; examples include Claude Desktop, Cursor, VS Code with MCP integrations, and agentic frameworks such as LangChain.
V-C Frameworks for Agentic Systems
Several frameworks provide abstractions for building complex, multi-step, and multi-agent workflows. These frameworks vary in their focus (retrieval, orchestration, or collaboration) but share the goal of simplifying agent construction and deployment. In this section, we explore key aspects and highlight the use cases of several prominent frameworks, with more detailed attention to the LangChain and on top of that LangGraph framework, which has gained notable popularity in recent agentic robotics systems.
V-C1 LangGraph/LangChain
LangGraph is a graph-based framework (nodes and edges) built on top of LangChain. To situate it briefly, LangChain is a modular framework that provides structured components for building tool-integrated and memory-aware agents based on LLMs. While LangChain focuses on modularity, LangGraph introduces a state-machine paradigm and supports conditional, non-linear workflows such as loops and retrievals. This makes it particularly suitable for building robust agent pipelines and even coordinating multiple interacting agents.
In LangGraph, each node represents a discrete piece of logic—this can be a call to an LLM, a function for data processing, or even another agent—while the edges define the control flow, including conditional branching and iterative loops. Several features are especially relevant for designing agentic physical AI systems:
  • Pre-built implementations: LangGraph provides ready-to-use templates for common agent patterns. Examples include:
    • Single-tool calling ReAct: Implements the ReAct pattern, combining reasoning traces with tool use in a structured way.
    • Supervisor multi-agent system: Individual agents are coordinated by a central supervisor agent, which controls all communication flow and task delegation. The supervisor decides which agent to invoke based on the current context and task requirements.
    • Swarm multi-agent system: Agents dynamically hand off control to one another depending on their specializations. The system tracks which agent was last active so that subsequent interactions resume with the same agent, ensuring conversational and task continuity.
  • MCP adapters: The framework integrates with the Model Context Protocol (MCP), making it easier to connect external tools and services in a standardized way.
  • Human-in-the-Loop capabilities: Developers can insert checkpoints where human users review or approve agent outputs before execution continues, ensuring safety and reliability.
  • Memory: LangGraph natively supports short- and long-term memory, enabling agents to retain context across steps or sessions, which is critical for persistent autonomy in physical environments.
V-C2 LlamaIndex
LlamaIndex specializes in RAG, providing structured ingestion and query pipelines over heterogeneous data sources. For agentic physical AI, this allows agents to ground decisions in sensor logs, spatial maps, or multimodal datasets. It is frequently paired with orchestration frameworks such as LangGraph for more complex workflows.
V-C3 CrewAI
CrewAI focuses on multi-agent collaboration, where agents are assigned roles and tasks within a coordinated “crew.” This makes it suitable for collaborative problem solving, distributed control, and simulation.
V-C4 AutoGen
AutoGen (Microsoft) emphasizes conversational multi-agent systems, supporting both AI–AI and AI–human collaboration. Its modular design makes it extensible to embodied AI scenarios, where agents must interact with humans and physical systems in tandem.
V-D Domain-specific considerations
Robotic data and control pipelines present unique challenges for agentic integration due to their inherent multimodality and temporal diversity. Typical robotic systems combine heterogeneous data streams (vision, laser ranging, tactile sensing, audio, joint encoders, or radio telemetry) each with distinct payload sizes, update frequencies, and latency tolerances. High-bandwidth modalities such as camera or depth feeds require continuous streaming and compression strategies, while others, like status or event messages, demand reliability and strict quality-of-service guarantees. Designing an agentic layer that can flexibly operate across these heterogeneous data rates and protocols remains nontrivial. Agents must reason over asynchronous information and decide which data to process, summarize, or ignore in real time, often under constraints of compute, bandwidth, and safety.
Another domain-specific constraint lies in embodiment variability. A general-purpose agent should be capable of interfacing with multiple robotic platforms; even if constrained to mobile agents, at least: mobile bases, mobile manipulators, drones, or legged robots (e.g., quadrupeds or humanoids). Each of these embodiments comes with different control interfaces and sensing configurations. Middlewares such as ROS or ROS 2 already provide abstraction layers, from abstracting hardware drivers with topics, services, and actions, to providing out-of-the-box manipulation (e.g., moveit) or navigation (e.g., nav2) capabilities. However, the degree to which an LLM- or VLM-driven agent can remain at a high level of reasoning depends on the embodiment. While protocol-based and interface-level integrations can remain platform-agnostic, models that generate low-level actions or policies (e.g., VLA or RL-based systems) still face challenges in transferring across morphologies. Generalizing such embodied control requires richer world models and modular skill representations that can decouple reasoning from actuation.
Finally, many real-world robotic tasks are inherently long-horizon and involve unpredictable interaction with the physical environment. Unlike static API calls, physical actions introduce variable execution times, uncertainties, and potential failure modes. Tasks such as “navigate to gate A42” in an airport, or “deliver this object to the suite in the top floor”, involve navigation through dynamic, human-centric spaces, possibly requiring waiting, replanning, or the use of elevators and constrained corridors. For an agentic autonomy solution, reasoning about such extended temporal dependencies (potentially spanning minutes or hours) requires continuous situational awareness, progress monitoring, and adaptive recovery strategies. These factors underscore that agentic systems in robotics must not only manage data and tools, but also reason about time, uncertainty, and embodiment, bridging symbolic decision-making with the continuous realities of physical interaction.
VI Discussion
Agentic Embodied AI is a rapidly developing field where research is arguably in its infancy. While embedded policies such as VLAs or BLMs are relatively mature, the approaches to integrating foundation models into existing or new autonomy stacks are still underexplored. Below, we outline some of the key open questions and promising directions for future work.
VI-A AI Agents, AI Workflows and Agentic AI
The deployment of LLMs across systems with differing levels of autonomy, decision logic, and task scope necessitates a conceptual distinction between AI Workflows, AI Agents, and Agentic AI Systems. These categories differ primarily in how control, reasoning, and coordination are distributed between humans, models, and the surrounding infrastructure.
AI Workflows operate along predefined, human-engineered paths in which the sequence of tool use and model calls is explicitly programmed. Here, the LLM acts as a component within a deterministic pipeline, executing a bounded subtask such as summarization. This approach prioritizes predictability, control, and auditability, making it well-suited for regulated environments or applications requiring reproducible outputs [anthropic2024building].
In contrast, AI Agents are systems that delegate control of task execution to the model itself. Rather than following a fixed sequence, the LLM dynamically decides which actions to take: planning, executing, and iterating toward a defined goal. The agent perceives its environment, evaluates progress, and selects subsequent actions such as invoking tools. This autonomy enables agents to tackle open-ended problems that cannot be easily pre-specified in a workflow. However, this flexibility comes at the expense of predictability, interpretability, and verifiability  [anthropic2024building, sapkota2505ai].
Finally, Agentic AI Systems represent a higher level of organization, in which multiple autonomous agents collaborate, coordinate, and share memory to achieve complex or long-horizon objectives. Rather than focusing on individual problem-solving capabilities, the emphasis shifts to the emergent behaviors that arise from interaction, negotiation, and collective reasoning among agents. This conceptual progression, from isolated task execution toward emergent multi-agent organization, is increasingly discussed in the context of robotics and autonomous systems [sapkota2505ai].
VI-B Embodiment
Embodiment can be understood as the capacity of a cyber-physical system to process information about its physical component. It can be viewed as a spectrum: some systems possess limited awareness of their physical properties, while others exhibit more advanced self-modeling. At the simplest level, this involves reasoning about the system’s own physical capabilities: its actuator modalities or “skills,” as outlined in the Internet of Skills [guzman2022ios]. On the other hand, it has been argued that true embodiment requires a more comprehensive internal model. Duffy and Joue [duffy2000embodiment] introduced the notion of strong embodiment, describing agents that not only understand their actuator modalities but also use this understanding to explore their environment, reason about it, and infer its physical laws.
Therefore, embodiment can be understood not only as an agent’s capacity to reason about its own body, but also about the environment in which it operates. It is thus better viewed not as an intrinsic property of the agent, but as a trait emerging from the dynamic coupling between the agent and its environment. Consequently, an agent’s embodiment may be context-dependent: a robot designed to perceive and act within a structured warehouse may lose embodiment when deployed in an open field.
For this reason, in this work we adopt the broadest possible definition of embodiment when conducting our survey. All systems considered possess some capacity to model and reason about their physical component within their operational context. While this capacity differs in form and degree (ranging from direct sensorimotor feedback to abstract physical modeling) each system can nonetheless be regarded as embodied. This diversity shows that embodiment is not a binary property but a spectrum shaped by how agents process information within, and about their environments.
VI-C Applications to Multi-Robot/agent Systems
In this work, we focus on embodied agentic AI solutions, in contrast to other surveys that explicitly categorize into single-agent and multi-agent approaches [feng2025multi]. The term multi-agent does not necessarily imply multi-robot; its interpretation varies across frameworks. For instance, in systems such as RAI [rachwal2025rai], a single robot may host multiple agents, whereas in orchestration architectures, one agent may control multiple robots. Furthermore, many frameworks adopt a one-to-one mapping between robots and agents [kannan2024smart]. Nonetheless, understanding the distribution of agents among robots remains an important aspect, as it reintroduces the debate between centralized and decentralized coordination architectures.
A key challenge is the development of general multi-robot/agent embodied AI frameworks. While recent progress has largely targeted constrained tasks, current methods struggle with scalability, heterogeneous agent teams, and variable objectives. Extending the principles of single-agent generalist models to multi-agent settings introduces difficulties such as non-stationarity, intricate interactions, and multiple equilibria. Unlike the closed and static scenarios often assumed in research, real-world deployments involve dynamic, uncertain, and adversarial conditions with noisy observations, shifting tasks, and fluid team compositions. Achieving robust performance in such environments requires resilient policies, continual learning, and context-aware reasoning to adapt rapidly to novel agents and evolving tasks.
Finally, applications across domains highlight both the promise and the challenges of embodied multi-robot/agent AI. Robotics for manufacturing, logistics, and autonomous driving demand scalable coordination and adaptability; education and healthcare require socially appropriate, ethical, and privacy-preserving interactions; and defense applications necessitate robustness under adversarial conditions with human oversight. Beyond these, smart cities, simulations, and energy systems call for efficient, trustworthy, and safe agent collectives. Addressing these diverse needs will require progress in multimodal perception, collaborative and social learning, and ethical governance, ensuring that multi-robot embodied AI evolves into a reliable, interpretable, and impactful technology for society.
VI-D Ethical consideration
Ethical issues arise when considering the deployment of embodied agents with humans: the possibility of physical harm, misinformation, or privacy violations. These are compounded by social threats posed by such systems, such as the rapidly changing labor market, and lack of sufficient policies regarding responsibility [perlo2025embodied]. They are said to fail to address the division of obligations between manufacturers of such systems, model providers, and organizations deploying them in the real world. Addressing these distinctions is crucial for responsible development and application of embodied AI systems. However, it comes with a complex set of challenges.
It could be argued that introducing legislation forcing agentic AI system providers to comply with regulation could lead to a decrease in competition by making it harder for new companies to join the market [Mueller2021AIActCost] [UKGov2023AIGovernanceImpact]. Finding a balance between appropriate regulation, while not stifling innovation, is crucial for successful and ethical deployment of AI-driven cyber-physical systems.
VI-E AI Agents in the Edge-Cloud Continuum
The increasing distribution of computational intelligence across robots, edge nodes, and cloud infrastructures introduces new challenges for embodied agentic AI. Robotic agents often operate in dynamic conditions where compute, storage, and connectivity must be balanced across heterogeneous devices and administrative domains. Agentic systems must therefore reason about where and how to execute perception, reasoning, and planning processes, dynamically orchestrating components in response to context and performance requirements. Distributed and federated approaches have begun to address this by enabling semantic task decomposition, runtime component migration, and adaptive resource selection across the compute continuum [Acharya2025, Filinis2026, Baccour2022]. These methods emphasize low-latency execution for perception at the edge while offloading heavy reasoning or model adaptation to the cloud, supporting a resilient and scalable paradigm for embodied autonomy. As such, the edge–cloud continuum becomes not merely an infrastructure layer but an integral part of the embodied agent’s cognitive architecture.
VI-F Memory
Long-term and contextually grounded memory remains one of the least mature capabilities in current agentic architectures. Although tool-based integration servers, such as Bagel [bagel] and rosbag MCP, provide standardized access to past sensor data and execution traces [ros2mcp2025], most existing agents still operate over transient, session-bound context windows. The lack of persistent, semantically indexed memory prevents agents from maintaining situational continuity, sharing experience between tasks, or reasoning over past states in real time. Emerging frameworks explore structured and distributed memory abstractions to support multi-session persistence and retrieval across agents [Ehtesham2025, Mei2025], yet real-time synchronization and consistency remain open problems. Overcoming these limitations will require live, queryable memory layers capable of bridging symbolic reasoning with embodied experience—enabling agents to continuously learn, adapt, and recontextualize actions across time and embodiment.
VI-G Continuous data streams
Integrating data streams with LLMs, VLMs, and Embodied AI systems remains an active area of research.
In robotics, a central challenge lies in handling continuous data streams from sensors such as cameras, LiDAR, and radar. These streams provide real-time sensor input, but conflict with the discrete mode of operation of most LLMs and VLMs. Continuous data must therefore be transformed into suitable representations. For example, a 10 Hz video feed cannot be directly used by an LLM which takes more than 0.1 seconds to process an image. That would lead to desynchronization and potential information loss.
Recent work explores strategies for selecting and summarizing data streams. For example, Hu [hu2025mllmbasedvideoframe] introduced a method using a multimodal LLM for the selection of frames from a video. Similarly, Lin [lin2025captionskeyframeskeyscoremultimodal] proposed a keyframe identification model that uses captioning and scoring to preserve context. Such approaches highlight promising directions for aligning continuous perception with discrete reasoning. However, integrating streaming data into embodied AI systems remains an open challenge.
VI-H World Models
The concept of world models has recently gained significant attention in the field of robotics and artificial intelligence. World models are internal representations that an agent (e.g., a robot) constructs to reason about its environment, predict the consequences of its actions, and plan its behavior. One of the pioneering works on world models is the paper by Ha and Schmidhuber[world_models_2018], which proposed a framework for building compact, compressed representations of the world using deep neural networks. The authors demonstrated the effectiveness of world models in various reinforcement learning tasks, where the agent learns to navigate and complete tasks in simulated environments. Following the work of Ha and Schmidhuber, several researchers have explored different aspects of world models and their applications in robotics. For a recent survey please see [Kawaharazuka_2024]. Recent works are for instance [hafner2024masteringdiversedomainsworld], proposing DreamerV3, a general algorithm that outperforms specialized methods across over 150 diverse tasks, or [hansen2024tdmpc2scalablerobustworld], introducing TD-MPC2, a model-based RL algorithm designed for learning generalist world models on large non-curated datasets composed of multiple task domains, embodiments, and action spaces. The authors of [zhou2024dinowmworldmodelspretrained] present DINO-WM, a method for training visual models by using pretrained DINOv2 embeddings of image frames: once trained, given a target observation, one can directly optimize agent behavior by planning through DINO-WM using model-predictive control (MPC).
World models can be valuable tools to enhance the robot’s understanding of the environment, more specifically to gauge the consequences/effects of its actions on the environment. Coupled with an exploratory strategy (e.g., an agentic loop) to assess what possible actions could be undertaken to perform a specific task and their predicted effects, the robot can choose to enact the most promising strategies from the world model in the real world, leading to safer and more robust task execution. While a world model is learnt and possibly updated from ”experience”, another similar exploration approach could be implemented by using simulation.
VI-I Evaluation Metrics
Evaluation metrics for embodied agentic systems can be defined along the Perceive–Reason–Act paradigm. Task performance is often measured through success rate and completion time, particularly in multi-stage navigation and manipulation tasks. The efficiency of tool and function definition significantly influences these outcomes, as clear and adaptable tool descriptions improve task execution regardless of the underlying tool-use capabilities of the foundation models. A critical aspect of evaluating physical agentic systems is the level of autonomy, defined as the ability of a system to plan, decide, and act without constant human intervention. Many existing frameworks demonstrate a medium level of autonomy, as they rely on predefined commands, functions, or tool sets. Achieving higher autonomy requires agents to generate their own sub-goals and corresponding actions. In some settings, especially those involving limited human interaction, this trait may be desirable. However, evaluating systems with higher autonomy is much more difficult.
Systems with a low degree of autonomy, i.e., those which utilize advanced tooling, can be often evaluated through monitoring the order and arguments of the tools. The higher the degree of autonomy, the need arises to evaluate the system through its impact on the environment in which it operates. This can lead to some undesirable behaviors, such as excessive movement, being missed during the evaluation. For example, in a highly autonomous agent responsible for its own path planning, simply checking whether it reached the goal may overlook inefficient or excessive movement. That issue is not present in low-autonomy systems using predefined navigation tools.
Adaptability is another key factor, reflecting the agent’s ability to generalize across unseen tasks, environments, or robotic embodiments with minimal additional data or modification. Memory consistency and recallability assess how effectively an agent stores, retains, and reuses contextually relevant information, serving as a foundation for stable, context-aware behavior. Other factors, such as safety, transparency, are also essential in evaluating embodied AI agents [raptis2025agentic] and have been extensively discussed in recent surveys on agentic AI [bandi2025rise].
VII Conclusion